OpenStudy (anonymous):

How to do 6x+5y=31;x= -4 It says solve the equation for y then find the value of y for the given value x

OpenStudy (ivettef365):

just substitute the x for -4 so find what is the value of y so you have 6(-4) + 5y = 31 solve for y

OpenStudy (ivettef365):

do you know how to solve it ?

OpenStudy (anonymous):

\[-24-5y=31 /+24\]

OpenStudy (anonymous):

\[-5y=55 /\div-5\]

OpenStudy (anonymous):

y=-11

OpenStudy (ivettef365):

you have -24 + 5y = 31 then add 24 on both sides and you have 5y = 31 + 24 5y = 55

OpenStudy (ivettef365):

where did you get -5y ?

OpenStudy (ivettef365):

@L-Lawliet-L ??

OpenStudy (anonymous):

i divided 55 and 5 with 5

OpenStudy (anonymous):

i thought it was an other sign too, i thought it was substract.

OpenStudy (anonymous):

the right answer is 11

OpenStudy (anonymous):

Thanks guys, in my boo it says y=-6/5x plus 31/5 then the answer 11 ?,5 A

OpenStudy (ivettef365):

yes 11 is the correct answer